OpenStudy (anonymous):

In how many ways can five married couples attending a concert be seated in a row of ten seats if the members of each sex are seated together?

OpenStudy (anonymous):

these are a pain and what does "married couples" have to do with it?

OpenStudy (anonymous):

if the women are seated in the 5 seats on the left, there are \(5!\) ways to do that, similarly there are \(5!\) ways to seat the men on the right so for that it would be \(5!\times 5!\) but of course the men could be on the left and the women on the right so the actual answer is \[2\times 5!\times 5!\]

OpenStudy (anonymous):

that's right!!! you are so awesome!!!